MICQ(7) MICQ(7) mICQ-NOTE-THISFILEHASNOTYETBEENUPDATEDFORmICQ0.4.10. ,mICQ.mICQ. trans[|nr]... ,isonr.,,LANG,LCALLLCMESSAGES,!,autodefault.,,.,none unload.,,debug.,,fun. ˜ /.micq/i18n/*.i18n,"". ,,,117,"transde117default". tabs, save,˜ /.micq/micqrc,.,. r[].-.,-(.).,,(#). a[],.r. msguin|nickname[] uin|nickname.r. urluin|nicknameurlmessage ,urluin|nickname. smscellmessage SMScell,,+<><>. msga--.(.),(#). resenduin|nickname... uinnickname.,. uptime mICQ. last[contact] contact,uin,nick,. verbose[] .: 8. 16. 32. 64’. 256v6. 4096v7/v8. 8192v7/v8. 16284. 65536"--". 131072 "--". UK mICQ 1 MICQ(7) MICQ(7) 262144 . 2097152 ’"--". 4194304 ’/. ,-,-1-. clear. sound[] /. soundonline[] /,. soundoffline[] /,. qmICQ. auto[on|off] /.on,off-. autostatus status. autoaway[timeout||] -timeout,autoaway,,.-0.,. alter. auth[req|grant|deny|add]uin|nick uin|nick-,uin|nick,-uin|nick,-. rand. passicq-. :save,,˜ /.micq/micqrc(.micqrc(7)).,. search searchemail searchnick searchfirstlast email,nick(@)’.,. addnicknamealias adduin[nickname] uin-nickname,aliasnickname.nickname,uinnickname. :save,. remalias alias.alias,. :save,. togvisnickname ,nickname,. toginvisnickname ,nickname. UK mICQ 2 MICQ(7) MICQ(7) togignickname ,nickname. f[inger]uin|nickname infouin|nickname uin|nickname. e,-. w-. ee,-. ww-.. wide-. ewide,-. s[nickname] ,nickname,,. i. status[nickname] nickname.:IP,ICQ’,UIN,nick,. :,sww. reg. change[[]] .. online"". away[]

""..

na[]"".. occ[]"".. dnd[]"".. ffc[]"".. inv"". update (email,). other,-. about. setonoff.color,funnyquiet. peek.ICQ;. setr[].. peeruin|nick uinnick. open’--TCP. close’--. off’,. UK mICQ 3 MICQ(7) MICQ(7) auto. away"". na"". dnd"". occ"". ffc"". file. files[]... .’’,.’/’,,’.’,’. conn[nr] ’’nr. open’. login’. close’.’. remove ()’. select’.nr’UIN,’. contact[] -: show-. diff-(uin/nick),-. import --. micq(1),micqrc(5) JamesMorrison<ja2morrison@student.math.uwaterloo.ca>mICQ.ArtemSokolov <sokolov@ukr.net>. UK mICQ 4